# Digital Simulation of Non-Abelian Parafermions in Superconducting Circuits

### v3

@@ III Ground state preparation and creation of parafermions
Figure 2: Illustration of the ground state preparation and parafermion creation protocol. (a) The $\mathbb{Z}_{3}$ plaquette model with a $6\times 6$ array of qutrits. (b) Explicit procudure for the ground state preparation. The system is initialized with all qutrits in $\left|{0}\right\rangle$. Red dots indicate the representative qutrits used to prepare the light plaquettes, while green dots represent the control qutrits used to apply conditional two-qutrit gates for preparing the dark plaquettes. The plaquettes in the ground state are colored pale yellow for light plaquettes and gold for dark plaquettes. (c) Applying the generalized Hadamard gate to the representative qutrits prepares the light plaquettes in the ground state. (d) A circuit is applied to the dark plaquettes to prepare them in the ground state. (e) Circuits for preparing boundary dark plauettes in the ground state. (f) Measuring two neighboring qutrits (encircled by dashed line) creates a pair of parafermions. (g) An alternative method for creating parafermions involves ground state preparation on a lattice with a hole, represented by a light blue region.
@@ III.1 Ground state preparation
We consider the $\mathbb{Z}_{3}$ plaquette model with $6\times 6$ qutrits [Fig. 2(a)]. The system is initially in the product state $\left|{0}\right\rangle^{\otimes 36}$. The ground state of the model is long-range entangled, which requires two-qutrit entangled operations for the preparation. This can be experimentally achieved through the tunable cross-Kerr interaction between two qutrits [Fig. 1(d)]. A specific configuration for this tunable coupling is shown in Fig. 1(e), where nearest-neighbor transmons are coupled via frequency-tunable couplers. Adjusting the coupler frequency modifies the strength of the cross-Kerr interaction, which, in combination with single-qutrit operations, enables the realization of efficient conditional phase gates Luo et al. (2023). Alternatively, the differential AC Stark shift on two transmon qutrits with static coupling can also dynamically realize the conditional phase gate Goss et al. (2022).
To prepare the ground state $\left|{\Psi_{0}}\right\rangle$ in Eq. (2), we first select a set of representative qutrits on light plaquettes [red dots in Fig. 2(b)] and apply the generalized Hadamard gate to them. The generalized Hadamard gate is defined by the quantum Fourier transform Wang et al. (2020); Ye et al. (2011):
@@
which is the common eigenstate of $X$ and $X^{\dagger}$ with eigenvalue $1$. The generalized Hadamard gate can be implemented by decomposing the gate into time evolutions as $H^{(3)}=e^{-iH_{d}t}e^{-iH_{o}t}$, where $H_{o}=\sum_{i<j}m_{ij}\left|{i}\right\rangle\langle{j}|+h.c.$ is off-diagonal, and $H_{d}=\text{diag}(\phi_{0},\phi_{1},\phi_{2})$ is diagonal with complex parameters $m_{ij}$ and real parameters $\phi_{i}$. The generator $H_{o}$ is experimentally realized by simultaneously microwave driving transitions between the three pairs of energy levels of the qutrit, while $H_{d}$ is implemented by shifting the phases of drive fields in the resonant control pulse O’Brien et al. (2004).
After applying the $H^{(3)}$ to each representative qutrit, all the light plaquettes (including the light boundary plaquettes) are prepared in the ground state [Fig. 2(c)]. Next, we select the control qutrit on each dark bulk plaquette [green dots in Fig. 2(b)] and apply $CZ$, $CZ^{\dagger}$, and $CX$ gates to prepare all the dark plaquettes in the ground state [Figs. 2(d) and 2(e)]. The control two-qutrit gates must be applied layer by layer, ensuring that the state stored in the representative qudits remains unchanged until the controlled operations on their respective plaquettes are applied. Ground state preparation on the dark plaquettes does not cause any light plaquette to deviate from the ground state, as their stabilizers commute with each other. In this way, we prepare the entire system in the ground state. The depth of the ground state preparation circuit depends linearly on the size of the lattice ($\sqrt{n}$ in our set up, with $n$ the number of qutrits).
@@
Fidelity of the ground state preparation can be obtained by the standard quantum state tomography (QST) method to read out the density matrix Bianchetti et al. (2010). With the density matrix, the expectation value $\langle\Pi_{p}^{0}\rangle$ of the ground state projector on each plaquette can be computed. By repeating the experiment multiple times, one obtains the average value $\overline{\langle\Pi_{p}^{0}\rangle}$ for each plaquette. If $\overline{\langle\Pi_{p}^{0}\rangle}=1$, it indicates perfect preparation of the ground state on the plaquette $p$; if $\overline{\langle\Pi_{p}^{0}\rangle}<1$, it indicates that the plaquette deviates from the ground state. This is the typical method used to characterize ground state preparation in digital simulations with SC qubits, and it is extended to the qutrit case in this work.
@@ III.2 Creation of parafermions
An alternative method arises in the presence of natural holes in the lattice [Fig. 2(g)], which result from imperfections in the geometric arrangement of qutrits due to fabrication constraints in chip manufacturing. Exploiting this imperfection, a pair of parafermions can be created by preparing the ground state on the imperfect lattice. However, for the region containing the hole, additional control gates, such as the controlled-$XZ$ gate, are required to enforce the ground-state condition. An example featuring a minimal hole is illustrated in Fig. 2(g).
@@ IV FUSION OPERATIONs OF THE GENERATED PARAFERMIONS
In this section, we propose the scheme to realize the fusion of parafermions. The previous section presents a method for creating parafermion pairs through engineering dislocations. Nevertheless, since the parafermions remain adjacent, studying the fusion rules of an individual parafermion is difficult because the nearby parafermion can influence the fusion and introduce ambiguity in the interpretation of results. To address this, we generalize the code deformation method to the qudit setting, deriving the operators that facilitate the spatial movement of parafermions. This generalization is technically non-trivial due to the non-Hermitian nature of the stabilizers, requiring careful treatment of both the stabilizer and its complex conjugate. Physically, this property is intrinsically linked to the non-self-duality of anyons and the nonlinearity of parafermion operator transformations, a connection we leave for future work. The generalized code deformation method yields operators that enable the horizontal and vertical movement of parafermions along the lattice. This capability serves as a powerful tool for parafermion manipulation and further facilitates essential operations, such as braiding [see Sec. V]. Given its significance and fundamental role in this study, we first introduce the generalized code deformation method before combining it with string operators to establish a protocol for realizing the fusion of parafermions.
@@ IV.1 Generalized code deformation
Code deformation is a method used to convert one error-correcting code into another by modifying the stabilizers, enabling fault-tolerant logical gate operations within surface codes Vuillot et al. (2019); Bombin (2011). Beyond quantum information science, code deformation also finds applications in condensed matter physics. Topological surface codes, such as the toric code, are rooted in topological quantum field theory, which provides the framework for topological orders. In this context, code deformation involves altering the topological structure of the field theory, such as engineering topological defects, to enrich the properties of topological orders Lensky et al. (2023). While the method has been extensively studied in qubit systems, its application to qudit systems remains relatively unexplored. In this subsection, we generalize the code deformation method to qudit systems. This generalization enables the derivation of unitary operators that facilitate the spatial movement of parafermions across the lattice. Our work extends the applicability of code deformation to a broader class of qudit stabilizer codes.
@@
Figure 3: Generalized code deformation and the fusion of parafermions. (a) Code deformation that moves the parafermion vertically. (b) Code deformation that moves the parafermion horizontally. (c) Illustration of the creation of a composite excitation and braiding $e$ or $\overline{e}$ with one parafermion. After the braiding, the excitations fuse to the trivial topological charge. (d) Explicit procedure for realizing the fusion of parafermions. Observables are characterized by $\langle\Pi_{p^{\prime}}^{d-1}\rangle=\langle\Pi_{p}^{1}\rangle=1$. (e) Circuit and two-qutrit gates used in the procedure.
@@
Before proceeding to the realization of the operators $U_{h}$ and $U_{v}$ using quantum circuits, we briefly review the code deformation method used to move Majorana fermions in stabilizer codes with SC qubits Xu et al. (2023); Andersen et al. (2023) and compare it with our generalization. In that case, both the original stabilizer $S_{\text{old}}$ and the transformed stabilizer $S_{\text{new}}$ after moving the Majorana zero mode are Hermitian, simplifying the construction of a unitary operator that transforms the system into an eigenvalue-$1$ eigenstate of the new stabilizer. Specifically, since $S_{\text{old}}$ anticommutes with $S_{\text{new}}$, their product $iS_{\text{new}}S_{\text{old}}$ is Hermitian, motivating the ansatz of the unitary operator $e^{\theta S_{\text{new}}S_{\text{old}}}$, where the parameter $\theta=\frac{\pi}{4}$ is fixed by applying the operator to the eigenvalue-$1$ eigenstate of $S_{\text{old}}$ and requiring that the resulting state be the eigenvalue-$1$ eigenstate of $S_{\text{new}}$. Nevertheless, in the qudit case, this ansatz does not directly apply. The absence of Hermitian stabilizers complicates the construction of the unitary operator in Eq. (10), making the generalization significantly non-trivial.
@@
where $\theta_{1}=\frac{\pi}{4}$ and $\theta_{2}\approx 1.696\pi$. These rotation and diagonal matrices can be experimentally implemented using microwave drives for level transitions and phase shifts of drive pulses Yurtalan et al. (2020); Morvan et al. (2021). The circuit implementation for $U_{v}$ is not necessarily unique and optimization algorithms can enhance its efficiency. Recent advances in optimization techniques Srinivasan et al. (2024) have reduced the number of required gates, while machine learning methods offer further potential improvements. Once $U_{v}$ is implemented, the operator $U_{h}$ can be obtained using the circuit in Fig. 4(b), where the qutrit SWAP gate is implemented using the circuit in Fig. 4(c). Here, the $K^{(3)}$ gate is defined by $K^{(3)}\left|{j}\right\rangle=\left|{3-j}\right\rangle$.
Figure 4: (a) Circuit that realizes the $U_{v}$ operator. (b) Circuit that transforms the $U_{v}$ operator into the $U_{h}$ operator. (b) The qutrit SWAP circuit.
@@ IV.2 Realization of fusion operations
which is a special case of Eq. (4). This rule indicates that the parafermion provides the fusion channels that identify the $e\overline{m}$ and $\overline{e}m$ excitations with the trivial topological charge $1$. The related fusion rule is
@@ V Braiding of parafermion modes
After realizing the fusion of parafermions, we now turn to the braiding of them. In Sec. IV, we introduce the generalized code deformation method, which provides two unitary operators that move parafermions vertically and horizontally [Figs. 3(a) and 3(b)]. These operations enable the full braiding of parafermions. To characterize the braiding statistics, we define the computational space basis encoded by parafermions. We show that the computational space of $\mathbb{Z}_{3}$ parafermions naturally encodes logical qutrits rather than qubits. This distinction arises from the difference in fusion rules between parafermions and Majorana fermions, with the former having more fusion channels. Within this larger computational space, parafermions exhibit more intricate braiding statistics, making them harder to measure. We identify the expectation value of local ground-state projectors as an observable for characterizing braiding results.
To encode information, a minimum number of four parafermions is required [Fig. 5(a)]. In this configuration, there are only two parity operators, $\Lambda_{1}=\overline{\omega}\gamma_{1}\gamma_{2}^{\dagger}$ and $\Lambda_{3}=\overline{\omega}\gamma_{3}\gamma_{4}^{\dagger}$ with $\gamma$ the parafermion operator [see Appendix. B]. For convenience, the computational space is restricted to the two-dimensional subspace spanned by $\Lambda_{1}\Lambda_{3}=1$. With this restriction, four parafermions are sufficient to encode a logical qutrit. To find the basis of the computational space, we diagonalize the parity operator $\Lambda_{i}$ and obtain the eigenstates $\{\left|{0}\right\rangle_{\Lambda},\left|{1}\right\rangle_{\Lambda},\left|{2}\right\rangle_{\Lambda}\}$. The logical qutrit state is defined by:
@@
On the other hand, physically, the basis of the computational space is defined by the fusion tree of the four parafermions [Fig. 5(b)]. The specific transformation between the basis defined by the eigenstates of the parity operators and the fusion basis $\left|{e\overline{m}/\overline{e}m}\right\rangle_{\gamma_{1}\gamma_{2}}\otimes\left|{\overline{e}m/e\overline{m}}\right\rangle_{\gamma_{3}\gamma_{3}}$ requires selecting two base points on the lattice and constructing the parafermion operators using string operators whose start and end points are at these base points and satisfying the algebra in Eq. (24Bombin (2010); Xu et al. (2023). Nevertheless, in this work, we do not make this exact transformation by selecting base points. Instead, we take a more direct approach to measure the braiding consequences of the parafermions, which does not rely on the explicit relation between the fusion basis and the logical qutrit basis $\left|{j}\right\rangle_{\Lambda_{1}}\otimes\left|{3-j}\right\rangle_{\Lambda_{3}}$.
In the logical qutrit basis $\{\left|{\overline{j}}\right\rangle\}$, the half braiding operator $U_{2}$ that exchanges $\gamma_{2}$ and $\gamma_{3}$ is an off-diagonal matrix, exhibiting the non-Abelian statistics. In a specific representation, the matrix elements of $U_{2}$ are given by: $\langle{\overline{i}}|U_{2}\left|{\overline{j}}\right\rangle=\frac{1}{\sqrt{3}}c_{i-j},$ where $c_{j}=\omega^{\frac{j(j+5)}{2}}$ [see Appendix. B]. The full braiding matrix in the logical qutrit basis is:
@@
We now move on to the explicit procedure. In Fig. 5(d), we show the implementation of full braiding of parafermions. The process begins by creating four parafermion $\gamma_{1}$, $\gamma_{2}$, $\gamma_{3}$, and $\gamma_{4}$ from the ground state, arranged in the configuration shown in Fig. 5(a). In the ground state, measuring $\langle\Pi_{p_{i}}^{0}\rangle$ in these plaquettes yields $\langle\Pi_{p_{i}}^{0}\rangle=1$. We then move $\gamma_{3}$ to the bottom, followed by moving $\gamma_{2}$ to the right. After these steps, we move $\gamma_{2}$ back to its original position and finally return $\gamma_{3}$ to its starting point. After braiding, the parafermions $\gamma_{1}$ and $\gamma_{2}$ fuse to form a composite excitation, denoted by $f$, which consists of the components $1$, $\overline{e}m$, and $e\overline{m}$. Similarly, parafermions $\gamma_{3}$ and $\gamma_{4}$ fuse into another composite excitation $f^{\prime}$ [Fig. 5(c)]. In general, $f$ and $f^{\prime}$ are distinct, but they are related by an electric-magnetic transformation, where $e\leftrightarrow m$. According to Eq. (20), the probability of measuring the topologically trivial charge $1$ from the composite excitation $f$ or $f^{\prime}$ is always $\frac{1}{3}$. Therefore, when we measure the expectation value $\langle\Pi_{p_{i}}^{0}\rangle$ using the QST method on the plaquettes that support the dislocation, we obtain $\overline{\langle\Pi_{p_{i}}^{0}\rangle}=\frac{1}{3}$, averaged over multiple repeating experiments. This provides an experimental observable for the braiding of parafermions. In Fig. 5(e), we show the two-qutrit gates used in this procedure.

### v4

@@ III Ground state preparation and creation of parafermions
Figure 2: Illustration of the ground state preparation and parafermion creation protocol. (a) The $\mathbb{Z}_{3}$ plaquette model with a $6\times 6$ array of qutrits. (b) Explicit procudure for the ground state preparation. The system is initialized with all qutrits in $\left|{0}\right\rangle$. Red dots indicate the representative qutrits used to prepare the light plaquettes in the ground state, while green dots represent the control qutrits used to apply controlled two-qutrit gates for preparing the dark plaquettes in the ground state. The plaquettes in the ground state are colored pale yellow for light plaquettes and gold for dark plaquettes. (c) Applying the generalized Hadamard gate to the representative qutrits prepares the light plaquettes in the ground state. (d) A circuit is applied to the dark plaquettes to prepare them in the ground state. (e) Circuits for preparing boundary dark plauettes in the ground state. (f) Measuring two neighboring qutrits (encircled by dashed line) creates a pair of parafermions. (g) An alternative method for creating parafermions involves ground state preparation on a lattice with a hole, represented by a light blue region.
@@ III.1 Ground state preparation
We consider the $\mathbb{Z}_{3}$ plaquette model with $6\times 6$ qutrits [Fig. 2(a)]. The system is initially in the product state $\left|{0}\right\rangle^{\otimes 36}$. The ground state of the model is long-range entangled, which requires two-qutrit entangled operations for the preparation. This can be experimentally achieved through the tunable cross-Kerr interactions between two qutrits [Fig. 1(d)]. A specific configuration for this tunable coupling is shown in Fig. 1(e), where nearest-neighbor transmons are coupled via frequency-tunable couplers. Adjusting the frequency of the coupler modifies the strength of the cross-Kerr interaction, which, in combination with single-qutrit operations, enables the realization of efficient conditional phase gates [81]. Alternatively, the conditional phase gate can also be dynamically realized by the differential AC Stark shift on two transmon qutrits with static coupling [84].
To prepare the ground state $\left|{\Psi_{0}}\right\rangle$ in Eq. (2), we first select a set of representative qutrits on light plaquettes [red dots in Fig. 2(b)] and apply the generalized Hadamard gate to them. The generalized Hadamard gate is defined by the quantum Fourier transform [55, 89]:
@@
which is the common eigenstate of $X$ and $X^{\dagger}$ with eigenvalue $1$. Experimentally, the generalized Hadamard gate can be implemented by decomposing the gate into time evolutions as $H^{(3)}=e^{-iH_{d}t}e^{-iH_{o}t}$, where $H_{o}=\sum_{i<j}m_{ij}\left|{i}\right\rangle\langle{j}|+h.c.$ is off-diagonal, and $H_{d}=\text{diag}(\phi_{0},\phi_{1},\phi_{2})$ is diagonal with complex parameters $m_{ij}$ and real parameters $\phi_{i}$. The generator $H_{o}$ is experimentally realized by simultaneously microwave driving transitions between the three pairs of energy levels of the qutrit, while $H_{d}$ is implemented by shifting the phases of drive fields in the resonant control pulse [90].
After applying the $H^{(3)}$ to each representative qutrit, all the light plaquettes (including the light boundary plaquettes) are prepared in the ground state [Fig. 2(c)]. Next, we select the control qutrit on each dark bulk plaquette [green dots in Fig. 2(b)] and apply $CZ$, $CZ^{\dagger}$, and $CX$ gates to prepare all the dark plaquettes in the ground state [Figs. 2(d) and 2(e)]. The controlled two-qutrit gates must be applied layer by layer, ensuring that the state of the representative qutrit remains unchanged until the controlled operations on their respective plaquettes are applied. Ground state preparation on the dark plaquettes does not cause any light plaquette to deviate from the ground state, as their stabilizers commute with each other. In this way, we prepare the entire system in the ground state. The depth of the ground state preparation circuit depends linearly on the size of the lattice ($\sqrt{n}$ in our set up, with $n$ the number of qutrits).
@@
Fidelity of the ground state preparation can be obtained by the standard quantum state tomography (QST) method to read out the density matrix [91]. With the density matrix, the expectation value $\langle\Pi_{p}^{0}\rangle$ of the ground state projector on each plaquette can be computed. By repeating the experiment multiple times, one obtains the average value $\overline{\langle\Pi_{p}^{0}\rangle}$ for each plaquette. If $\overline{\langle\Pi_{p}^{0}\rangle}=1$, it indicates perfect preparation of the ground state on the plaquette $p$; if $\overline{\langle\Pi_{p}^{0}\rangle}<1$, it indicates that the plaquette deviates from the ground state. This is the typical method used to characterize ground state preparation in digital simulations with SC qubits, and it is extended to the qutrit case in this work.
@@ III.2 Creation of parafermions
An alternative method arises in the presence of natural holes in the lattice [Fig. 2(g)], which result from imperfections in the geometric arrangement of qutrits due to fabrication constraints in chip manufacturing. Exploiting this imperfection, a pair of parafermions can be created by preparing the ground state on the imperfect lattice. However, for the region containing the hole, additional controlled gates, such as the controlled-$XZ$ gate, are required to enforce the ground-state condition. An example featuring a minimal hole is illustrated in Fig. 2(g).
@@ IV FUSION OPERATIONs OF THE GENERATED PARAFERMIONS
In this section, we propose the scheme to realize the fusion of parafermions. The previous section presents a method for creating parafermion pairs through engineering dislocations. Nevertheless, since the parafermions remain adjacent, studying the fusion rules of an individual parafermion is difficult because the nearby parafermion can influence the fusion and introduce ambiguity in the interpretation of results. To address this, we generalize the code deformation method to the qudit setting, deriving the operators that facilitate the spatial movement of parafermions. This generalization is technically non-trivial due to the non-Hermitian nature of the stabilizers, requiring careful treatment of both the stabilizer and its Hermitian conjugate. Physically, this property is intrinsically linked to the non-self-duality of anyons and the nonlinearity of parafermion operator transformations, a connection we leave for future work. The generalized code deformation method yields operators that enable the horizontal and vertical movement of parafermions along the lattice. This capability serves as a powerful tool for parafermion manipulation and further facilitates essential operations, such as braiding [see Sec. V]. Given its significance and fundamental role in this study, we first introduce the generalized code deformation method before combining it with string operators to establish a protocol for realizing the fusion of parafermions.
@@ IV.1 Generalized code deformation
Code deformation is a method used to convert one error-correcting code into another by modifying the stabilizers, enabling fault-tolerant logical gate operations within surface codes [92, 93]. Beyond quantum information science, code deformation also finds applications in condensed matter physics. Topological surface codes, such as the toric code, are rooted in topological quantum field theory, which provides the framework for topological orders. In this context, code deformation involves altering the topological structure of the field theory, such as engineering topological defects, to enrich the properties of topological orders [94]. While the method has been extensively studied in qubit systems, its application to qudit systems remains relatively unexplored. In this subsection, we generalize the code deformation method to the qudit system. This generalization enables the derivation of unitary operators that facilitate the spatial movement of parafermions across the lattice. Our work extends the applicability of code deformation to a broader class of qudit surface codes.
@@
Figure 3: Generalized code deformation and the fusion of parafermions. (a) Code deformation that moves the parafermion vertically. (b) Code deformation that moves the parafermion horizontally. (c) Illustration of the creation of a composite excitation and braiding $e$ or $\overline{e}$ with one parafermion. After the braiding, the excitations fuse to the trivial topological charge. (d) Explicit procedure for realizing the fusion of parafermions. Observables are characterized by $\langle\Pi_{p^{\prime}}^{d-1}\rangle=\langle\Pi_{p}^{1}\rangle=1$. (e) Quantum circuit and two-qutrit gates used in the procedure.
@@
Before proceeding to the realization of the operators $U_{h}$ and $U_{v}$ using quantum circuits, we briefly review the code deformation method used to move Majorana modes in stabilizer codes with SC qubits [3, 4] and compare it with our generalization. In that case, both the original stabilizer $S_{\text{old}}$ and the transformed stabilizer $S_{\text{new}}$ after moving the Majorana mode are Hermitian, simplifying the construction of a unitary operator that transforms the system into an eigenvalue-$1$ eigenstate of the new stabilizer. Specifically, since $S_{\text{old}}$ anticommutes with $S_{\text{new}}$, their product $iS_{\text{new}}S_{\text{old}}$ is Hermitian, motivating the ansatz of the unitary operator $e^{\theta S_{\text{new}}S_{\text{old}}}$, where the parameter $\theta=\frac{\pi}{4}$ is fixed by applying the operator to the eigenvalue-$1$ eigenstate of $S_{\text{old}}$ and requiring that the resulting state be the eigenvalue-$1$ eigenstate of $S_{\text{new}}$. Nevertheless, in the qudit case, this ansatz does not apply. The absence of Hermitian stabilizers complicates the construction of the unitary operator in Eq. (10), making the generalization significantly non-trivial.
@@
where $\theta_{1}=\frac{\pi}{4}$ and $\theta_{2}\approx 1.696\pi$. These rotation and diagonal matrices can be experimentally implemented using microwave drives for level transitions and phase shifts of drive pulses [82, 83]. The circuit implementation for $U_{v}$ is not necessarily unique and optimization algorithms can enhance its efficiency [95]. Machine learning methods may offer further potential improvements. Once $U_{v}$ is implemented, the operator $U_{h}$ can be obtained using the circuit in Fig. 4(b), where the qutrit SWAP gate is implemented using the circuit in Fig. 4(c). Here, the $K^{(3)}$ gate is defined by $K^{(3)}\left|{j}\right\rangle=\left|{3-j}\right\rangle$.
Figure 4: (a) Quantum circuit that realizes the $U_{v}$ operator. (b) Quantum circuit that transforms the $U_{v}$ operator into the $U_{h}$ operator. (b) The qutrit SWAP circuit.
@@ IV.2 Realization of fusion operations
which is a special case of Eq. (4). This rule indicates that the parafermion provides the fusion channels that identify the $e\overline{m}$ and $\overline{e}m$ excitations with the trivial topological charge $1$. The related fusion rule is:
@@ V Braiding of parafermion modes
After realizing the fusion of parafermions, we now turn to the braiding of them. In Sec. IV, we introduce the generalized code deformation method, which provides two unitary operators that move parafermions vertically and horizontally [Figs. 3(a) and 3(b)]. These operations enable the full braiding of parafermions. To characterize the braiding statistics, we define the computational space basis encoded by parafermions. Compared to Majorana modes, the computational space of $\mathbb{Z}_{3}$ parafermions naturally encodes logical qutrits rather than qubits. This distinction arises from the difference in fusion rules between parafermions and Majorana mdoes, with the former having more fusion channels. Within this larger computational space, parafermions exhibit more intricate braiding statistics, making them harder to measure. We identify the expectation value of local ground-state projectors as an observable for characterizing braiding results.
To encode information, a minimum number of four parafermions is required [Fig. 5(a)]. In this configuration, there are only two parity operators, $\Lambda_{1}=\overline{\omega}\gamma_{1}\gamma_{2}^{\dagger}$ and $\Lambda_{3}=\overline{\omega}\gamma_{3}\gamma_{4}^{\dagger}$ with $\gamma$ the parafermion operator [see Appendix. B]. For convenience, the computational space is restricted to the two-dimensional subspace spanned by $\Lambda_{1}\Lambda_{3}=1$. With this restriction, four parafermions are sufficient to encode a logical qutrit. To find the basis of the computational space, we diagonalize the parity operator $\Lambda_{i}$ and obtain the eigenstates $\{\left|{0}\right\rangle_{\Lambda_{i}},\left|{1}\right\rangle_{\Lambda_{i}},\left|{2}\right\rangle_{\Lambda_{i}}\}$. The logical qutrit state is defined by:
@@
On the other hand, physically, the basis of the computational space is defined by the fusion tree of the four parafermions [Fig. 5(b)]. The specific transformation between the basis defined by the eigenstates of the parity operators and the fusion basis $\left|{\gamma_{1},\gamma_{2};1/e\overline{m}/\overline{e}m}\right\rangle\otimes\left|{\gamma_{3},\gamma_{4};1/\overline{e}m/e\overline{m}}\right\rangle$ requires selecting two base points on the lattice and constructing the parafermion operators using string operators whose start and end points are at these base points and satisfying the algebra in Eq. (24[18, 3]. Nevertheless, in this work, we do not make this exact transformation by selecting base points. Instead, we take a more direct approach to measure the braiding consequences of the parafermions, which does not rely on the explicit relation between the fusion basis and the logical qutrit basis $\left|{j}\right\rangle_{\Lambda_{1}}\otimes\left|{3-j}\right\rangle_{\Lambda_{3}}$.
In the logical qutrit basis $\{\left|{\overline{j}}\right\rangle\}$, the half-braiding operator $U_{2}$ that exchanges $\gamma_{2}$ and $\gamma_{3}$ is an off-diagonal matrix, exhibiting the non-Abelian statistics. In a specific representation, the matrix elements of $U_{2}$ are given by: $\langle{\overline{i}}|U_{2}\left|{\overline{j}}\right\rangle=\frac{1}{\sqrt{3}}c_{i-j},$ where $c_{j}=\omega^{\frac{j(j+5)}{2}}$ [see Appendix. B]. The full braiding matrix in the logical qutrit basis is:
@@
We now move on to the explicit procedure. In Fig. 5(d), we show the implementation of full braiding of parafermions. The process begins by creating four parafermion $\gamma_{1}$, $\gamma_{2}$, $\gamma_{3}$, and $\gamma_{4}$ from the ground state, arranged in the configuration shown in Fig. 5(a). In the ground state, measuring $\langle\Pi_{p_{i}}^{0}\rangle$ in these plaquettes yields $\langle\Pi_{p_{i}}^{0}\rangle=1$. We then move $\gamma_{3}$ to the bottom, followed by moving $\gamma_{2}$ to the right. After these steps, we move $\gamma_{2}$ back to its original position and finally return $\gamma_{3}$ to its starting point. After braiding, the parafermions $\gamma_{1}$ and $\gamma_{2}$ fuse to form a composite excitation, denoted by $f$, which consists of the components $1$, $\overline{e}m$, and $e\overline{m}$. Similarly, parafermions $\gamma_{3}$ and $\gamma_{4}$ fuse into another composite excitation $f^{\prime}$ [Fig. 5(c)]. In general, $f$ and $f^{\prime}$ are distinct, but they are related by an electric-magnetic transformation, where $e\leftrightarrow m$. According to Eq. (20), the probability of measuring the topologically trivial charge $1$ from the composite excitation $f$ or $f^{\prime}$ is always $\frac{1}{3}$. Therefore, when measuring the expectation value $\langle\Pi_{p_{i}}^{0}\rangle$ using the QST method on the plaquettes that support the dislocation, the averaged result over multiple repeated experiments is given by $\overline{\langle\Pi_{p_{i}}^{0}\rangle}=\frac{1}{3}$. This provides an experimental observable for the braiding of parafermions. In Fig. 5(e), we show the two-qutrit gates used in this procedure.